OpenStudy (chloed18):

the sides of the triangle at right are formed by the graphs of 3x + 2y = 1, y = x -2, and -4x + 9y = 22. Is the triangle isosceles? How do you know?

OpenStudy (anonymous):



OpenStudy (chloed18):

How? I don't have x and y

OpenStudy (chloed18):

@surjithayer ^

OpenStudy (anonymous):

3x+2y=1 y=x-2 3x+2(x-2)=1 3x+2x-4=1 5x=5 x=1 y=1-2=-1 so co-ordinates of one vertex say A are (1,-1) now consider -4x+9y=22 y=x-2 -4x+9(x-2)=22 -4x+9x=22+18 5x=40 x=8 y=8-2=6 so co-ordinates of vertex say C are (8,6) Again consider 3x+2y=1 -4x+9y=22